# A Comprehensive Survey on Machine Learning Driven Material Defect Detection: Challenges, Solutions, and Future Prospects

### v1

@@ 3.2.3. Ensemble Learning Algorithms
Ensemble learning in material defect detection involves using multiple learning algorithms or models to obtain better predictive performance than could be obtained from any of the individual models alone (alipour2020increasing, ). A novel method for robotic hammering inspection aimed at maintaining social infrastructures, with a specific focus on estimating material defects like delamination depth in concrete (fujii2016defect, ). With the aging of infrastructure, there’s a growing need for automated diagnostic methods. The hammering test, known for its efficiency in inspections, is gaining interest in robotic automation. The study introduces a method to estimate defect conditions suitable for detailed robotic inspections. It also incorporates an integration technique of multiple classifiers to enhance inspection accuracy. Additionally, the system is designed to minimize the impact of mechanical running noise. Experimental results with concrete test samples show the effectiveness of this approach, confirming its accuracy in both defect detection and condition estimation. In addition, a new deep ensemble learning method for real-time surface defect inspection in automated production focuses on overcoming challenges like domain shifts and imbalanced training data in flowing data scenarios (liu2019data, ). It introduces a unique distribution discrepancy identifier that blends deep CNNs with shallow learning for improved robustness and precision. The method was successfully tested on an industrial surface mount technology line, showing excellent mean average precision and adaptability for streaming data in industrial settings.
3.2.4. Attention Mechanisms Related Algorithms
Attention mechanisms in material defect detection refer to a set of techniques in machine learning, particularly in deep learning, that enable the model to focus on specific parts of the input data that are more relevant to the task at hand. Fang et al. (fang2022tactile, ) propose a novel fabric defect detection method using a vision-based tactile sensor, which is more robust than traditional visual methods as it’s unaffected by dyeing patterns and ambient light. A robotic arm with tactile sensors automates data collection, creating comprehensive fabric datasets. The approach employs a convolutional neural network (CNN) with an attention mechanism and frequency domain filtering to enhance defect detection accuracy and efficiency. Experiments demonstrate its superiority over visual defect detection methods, confirming its practicality and efficiency for real-world applications. In addition, Chen et al. (chen2021defect, ) present a method for detecting defects in aluminium profile surfaces (APS) using a deep self-attention mechanism (DSAM) specifically designed to work effectively in noisy image conditions. This method improves upon traditional approaches by incorporating a residual learning strategy for better feature extraction from noisy images, employing a self-attention mechanism to refine these features, and using a softmax classification layer to identify defects in APS accurately. The method has achieved higher accuracy in detecting APS defects under various noise conditions.
@@ 3.2.5. Supervised Technique Summary
This section summarizes the application of supervised learning techniques in material defect detection, including regression algorithms, neural network algorithms, ensemble learning algorithms, and attention mechanism-related algorithms, with a comparison provided in Table 3. Regression algorithms are suitable for predicting the severity and extent of defects, offering high detection accuracy, but requiring a large amount of labelled data and difficult adaptation to new types of defects. Neural network algorithms are applicable to complex and nonlinear problems, enabling accurate detection and quantitative assessment of defects, but demanding substantial training data and computational resources. Ensemble learning algorithms improve predictive performance by combining multiple models, enhancing detection accuracy and robustness, although the models may become overly complex and challenging to interpret. Attention mechanism-related algorithms enable models to focus on task-relevant parts, enhancing the accuracy and efficiency of defect detection, yet requiring adjustment and optimization to suit specific application scenarios. Considering task requirements, data characteristics, and computational resources, selecting the appropriate techniques is crucial for improving the accuracy and efficiency of material defect detection.
Table 3. Supervised Technique Comparison in Material Defect Detection
Supervised Technique Application Scenarios Advantages Limitations
Regression
Various types of material defects,
such as surface defects,
internal defects, etc.
accurate predictions of defectiveness
and severity, as well as quantitative
assessments of defects.
need Large amounts of labeled data
and be hard to adapt to new defects
Neural Network
Complex defect feature extraction and
classification with intricate
backgrounds, structures, and noise.
Strong adaptability to different
defect types, capable of learning
complex defect features and
accommodating large-scale
defect detection.
Requires extensive annotated defect
data, prone to model overfitting and
black-box decision-making.
Ensemble Learning
Multimodal defect types,
complex defect detection task
Enhanced defect detection performance
and model generalization, high flexibility
High algorithm complexity, high
hardware infrastructure requirements,
demanding defect data quality
Attention mechanisms
Detection of defects of
various types and scales
Enhanced defect detection accuracy
and efficiency by focusing on local
defect features, simultaneously
improving model adaptability
Increased computational complexity,
increased defect data volume and
annotation costs, poor model interpretability
@@ 3.3. Semi-supervised Learning Algorithms
In practical industrial scenarios, the collection and annotation costs of defective samples are often high, leading to a significant amount of unlabeled data. In order to make more effective use of these unlabeled data, researchers have turned to the setting of semi-supervised learning (liu2022defect, ; liu2021semisupervised, ; wang2021new, ). In semi-supervised learning, the training set includes both normal and abnormal samples, but only a small portion of the samples are labelled. Compared to supervised learning, semi-supervised learning reduces the dependence on labelled data.
@@
Compared to unsupervised methods, semi-supervised learning significantly improves model performance when provided with a small amount of labelled defective samples. This approach is beneficial in addressing new types of defects in industrial product manufacturing. The application of semi-supervised learning is primarily focused on defect classification at the image level, while research on defect localization tasks is still an area that requires further exploration.
@@ 3.4. Reinforcement Learning Algorithms
Material defect detection often faces complex and dynamic environments, where the types and features of defects may exhibit diversity. Reinforcement learning can learn optimal strategies through interaction with the environment, enhancing the effectiveness of material defect detection. Therefore, reinforcement learning, capable of meeting the demands of more complex material defect detection environments and achieving autonomous real-time dynamic learning of material defects, has started to gain favour among researchers.
@@
In (chu2020neural, ), a reinforcement learning-based neural batch sampler is employed to select training batches for an autoencoder that are more conducive to generating loss curves distinguishing abnormal and normal regions. Additionally, periodic reinitialization and retraining of the autoencoder allow it to start from different initial weights and optimize towards different local minima. Throughout the training process, reinforcement learning gradually reduces its dependence on behaviour cloning rewards while increasing reliance on prediction loss rewards, facilitating a smooth transition from a heuristic policy to a more optimized strategy. In addition, (chung2022reinforcement, ) introduces a reinforcement learning-based online defect detection and mitigation method to enhance quality assurance in additive manufacturing processes. The proposed algorithm, called Continuous G-Learning, simultaneously utilizes two types of prior knowledge (offline and online) to reduce the number of learning samples by adjusting process parameters, ensuring the learned policy is optimal. Experimental results demonstrate the ability of this method to rapidly detect and mitigate emerging defects, improving surface quality and mechanical performance of printed products. Also, In (chung2022reinforcement, ), a reinforcement learning-based virtual plastic texture surface defect generation system is proposed for data augmentation in deep learning detection systems. The reinforcement learning network, consisting of an actor network and a critic network, searches for optimal rendering and defect parameters to generate realistic texture images. Using the generated images to supplement the original training dataset enhances data diversity and balance, improving the detection system’s recall rate. Lastly, (zhang2023duak, ) collects and integrates heterogeneous steel surface defect data from a human-network-physical system, constructing a steel surface defect knowledge graph. The proposed bidirectional reinforcement learning-based knowledge graph inference method employs two agents exploring paths from opposite directions, leveraging the historical experience and path-sharing mechanisms to update action selection policies and reward functions, enhancing inference efficiency and accuracy. Applied to steel surface defect detection and diagnosis, this method leverages knowledge graph inference to supplement missing knowledge, providing decision support for process improvement. In (mohanty2024enhancing, ), a digital twin model of a titanium alloy washer is created using Blender and Blendtorch, incorporating geometry, surface texture, lighting conditions, and camera parameters. A Proximal Policy Optimization (PPO) algorithm is employed for reinforcement learning to train an agent, interacting with the Blender environment to learn how to adjust rendering parameters to deceive the discriminator and enhance authenticity. A defect detector based on ResNet18 is trained using synthetic defective and non-defective washer images and evaluated for performance and robustness on real washer images.
@@
Reinforcement learning methods can adapt to continuously changing environments in material defect detection. This is because they learn through interaction with the environment, enabling them to continuously adjust their strategies during training to accommodate new defect types or environmental changes. They also exhibit high real-time capability, allowing learning and decision-making in real-time data streams, which is crucial for scenarios requiring timely responses in material defect detection. Additionally, reinforcement learning models can autonomously learn defect features without the need for extensive manual feature engineering. This enables them to extract valuable information from data and exhibit adaptability to complex defect situations. However, reinforcement learning typically demands substantial training and computational resources, especially when dealing with complex material defect detection tasks. This may result in long training times and high computational costs. Moreover, reinforcement learning models usually require a large number of samples for training, particularly in the case of deep reinforcement learning models. Acquiring large-scale annotated data in the field of material defect detection can be an expensive and challenging task. Overall, reinforcement learning holds potential in material defect detection.
@@ 3.5. Generative Learning Algorithms
To obtain more diverse augmentation data, generative learning techniques are introduced into the augmentation of defect sample data, aiming to synthetically create realistic defects on normal samples. The work by (valente2020print, ) introduces an end-to-end deep learning framework based on semantic segmentation for mapping printing defects. This framework enables pixel-level detection and localization of defects, such as stripes and colour bands, in printed images. Two types of printing defects and printing scan effects are simulated using image processing and computer graphics techniques, generating synthetic training datasets. This approach avoids the costs and time associated with manually collecting and annotating data. Experimental results demonstrate that the proposed framework can effectively learn features from synthetic data and generalize them to real printing defects. In (bosse2023automated, ), a data-driven feature labelling model is proposed for detecting hidden defects, such as pores in aluminium alloy die-cast plates, in low-cost X-ray projection images. Due to the absence of accurate annotated data for real X-ray images, the authors use numerical simulations of X-ray images to generate synthetic training datasets, taking into account various noise and distortion factors. Furthermore, (farady2023preaugnet, ) introduces an offline PreAugmentation Network (PreAugNet) for generating and filtering augmented samples to improve the training of target networks. The PreAugmentation Network utilizes a convolutional neural network (CNN) to extract features from original samples and updates the generated samples through the decision boundary analysis of a support vector machine (SVM) classifier. This ensures that the new samples produced by the PreAugmentation Network are more diverse and representative. Evaluation on three real industrial steel surface defect datasets demonstrates the effectiveness of PreAugmentation Network data augmentation in improving the training of defect classification networks.
@@
In addition to the methods mentioned above for generating defect data, the most common generative techniques are autoencoders and Generative Adversarial Networks (GANs). (yun2020automated, ) proposes a metal surface defect classification algorithm based on Deep Convolutional Neural Networks (DCNN), utilizing metal surface images collected by a machine vision system as input. To address the issue of imbalanced defect data, they introduce a data augmentation approach based on the Conditional Convolutional Variational Autoencoder (CCVAE). By learning the distribution of defect data, CCVAE generates images of various defect classes. Training the DCNN classification model on data generated by CCVAE enhances the model’s generalization performance and accuracy. Besides autoencoders, GANs is the most popular technique employed to generate the defect images. (pourreza2021g2d, ) presents a defect detection method based on Generative Adversarial Networks (GAN), named G2D (Generate to Detect Anomaly), for detecting anomalies in images and videos with only normal samples. G2D leverages the generator of GAN to produce different types of anomaly samples during training, including randomly generated samples deviating from the normal distribution and semantically inconsistent samples close to the boundary. (zhang2021defect, ) introduces Defect-GAN, a defect sample synthesis method based on Generative Adversarial Networks (GAN), to address the issue of insufficient defect samples in automatic defect detection. Defect-GAN simulates the damage and restoration process on a large number of normal samples to generate realistic and diverse defect samples. (jain2022synthetic, ) proposes a data augmentation method based on Generative Adversarial Networks (GAN) to improve the performance of surface defect detection and classification. Three different GAN architectures (DCGAN, AC-GAN, and InfoGAN) are employed to generate surface defect images containing six types of hot-rolled steel strip defects. Additionally, the study explores the latent space of GANs, discovering that adjusting continuous or discrete latent variables can control certain features of synthetic images, such as the severity of defects and lighting conditions, thereby incorporating additional domain knowledge.
@@
In summary, generative learning methods can be employed in material defect detection to synthesize defect samples, thereby increasing the diversity of training data. This is particularly beneficial for material defect detection tasks that lack a large number of annotated defect samples, as it can significantly alleviate the cost of defect annotation, address data imbalance issues, and enhance model robustness. However, the performance of generative learning methods largely depends on the quality of the generated model, which may be constrained by the limitations of training data and may struggle to fully replicate the real-world distribution of defects. Additionally, synthesized defect samples may not be sufficiently diverse, leading to a decline in model performance in practical applications. Moreover, some generative learning methods, especially those based on deep learning, may have high computational resource requirements, especially when training and generating on a large scale. Therefore, the choice of generative learning methods should be comprehensively considered based on specific task requirements, data characteristics, and computational resources.

### v3

@@ 3.2.3. Ensemble Learning-Based Techniques
Ensemble learning in MDD involves using multiple learning algorithms or models to obtain better predictive performance than could be obtained from any of the individual models alone (alipour2020increasing, ). A novel method for robotic hammering inspection aimed at maintaining social infrastructures, with a specific focus on estimating material defects like delamination depth in concrete (fujii2016defect, ). With the aging of infrastructure, there’s a growing need for automated diagnostic methods. The hammering test, known for its efficiency in inspections, is gaining interest in robotic automation. The study introduces a method to estimate defect conditions suitable for detailed robotic inspections. It also incorporates an integration technique of multiple classifiers to enhance inspection accuracy. Additionally, the system is designed to minimize the impact of mechanical running noise. Experimental results with concrete test samples show the effectiveness of this approach, confirming its accuracy in both defect detection and condition estimation. In addition, a new deep ensemble learning method for real-time surface defect inspection in automated production focuses on overcoming challenges like domain shifts and imbalanced training data in flowing data scenarios (liu2019data, ). It introduces a unique distribution discrepancy identifier that blends deep CNNs with shallow learning for improved robustness and precision. The method was successfully tested on an industrial surface mount technology line, showing excellent mean average precision and adaptability for streaming data in industrial settings.
Table 5. Comparison of supervised learning techniques for material defect detection
Technique Category Application Scenarios Advantages Limitations
Regression-Based
Quantitative defect
characterization (e.g.,
depth, size, severity)
Provides precise numerical
predictions and continuous
outcome estimation
Highly dependent on labeled data
and limited generalizability
for unseen defect types
Neural Network
Diverse defect types
and complex,
nonlinear problems
Adaptable to varied defect
scenarios; learns intricate
defect patterns
Requires large annotated
datasets; computational
overhead
- Convolutional Neural Network
Image-based defect
detection tasks
Robust texture analysis,
scalable to large datasets
Sensitive to hyperparameter
selection; computationally
expensive
- Recurrent Neural Network
Time-series analysis
in defect detection
Captures temporal patterns
for predictive maintenance
Prone to vanishing gradient
issues; less suitable
for large-scale data
- Transformer Models
Global feature analysis
for high-resolution
defect detection
Captures global context;
excels in processing
structured data
Requires high computational
resources; needs large
datasets for training
- Graph Neural Networks
Relational defect
modeling (e.g.,
microstructures)
Models complex relationships
in structured data
Limited scalability; sensitive
to graph quality
- Hybrid Neural Network
Multimodal defect
detection combining
data types
Combines multiple
architectures for improved
performance
Increased model complexity
and resource requirements
Attention Mechanisms
Localized defect detection
in complex environments
Enhances focus on relevant
features, boosting accuracy
Requires substantial computational
power; optimization
challenges
Ensemble Learning
Combining models for
complex defect scenarios
Improved robustness;
enhanced model
generalization
High implementation complexity;
demanding on data
quality
@@ 3.2.4. Supervised Technique Summary
This section summarizes the application of SL techniques in MDD, including regression-based techniques, neural network-based algorithms, and ensemble learning techniques, with a comparison provided in Table 5. Regression-based techniques provide precise numerical predictions, making them suitable for quantitative defect characterization, such as depth, size, or severity, but they require highly annotated data and exhibit limited generalizability to unseen defect types. Neural network algorithms are applicable to complex and nonlinear problems, enabling accurate detection and quantitative assessment of defects, but they are further categorized into subtypes such as CNNs, RNNs, Transformer models, GNNs, and Hybrid Neural Networks, each offering unique advantages tailored to specific defect detection scenarios. For instance, CNNs excel in image-based tasks, while GNNs model structured relational data. These approaches demand extensive annotated datasets and incur significant computational costs. Attention mechanisms enable models to focus on task-relevant parts, enhancing the accuracy and efficiency of defect detection, particularly in localized and complex environments, but their integration increases computational and optimization complexity. Ensemble learning techniques improve predictive performance by combining multiple models, enhancing detection accuracy and robustness, although their high implementation complexity and reliance on high-quality defect data may pose challenges in practical deployment. Considering task requirements, data characteristics, and computational resources, selecting the appropriate techniques is crucial for improving the accuracy and efficiency of MDD.
@@ 3.3. Semi-supervised Learning Algorithms
In practical industrial scenarios, the collection and annotation costs of defective samples are often high, leading to a significant amount of unlabeled data. In order to make more effective use of these unlabeled data, researchers have turned to the setting of SSL (liu2022defect, ; liu2021semisupervised; wang2021new, ). In SSL, the training set includes both normal and abnormal samples, but only a small portion of the samples are labeled. Compared to SL, SSL reduces the dependence on labeled data.
@@
Similar to (he2019semi, ), (gao2020semi, ) introduces an SSL approach based on CNN for steel surface defect identification. It leverages a small set of labeled samples and a large amount of unlabeled samples for model training. The model’s generalization capability is enhanced by generating pseudo-labels for unlabeled samples, thereby using the pseudo-labeling technique. Also, the framework proposed by (dai2020soldering, ) encompasses a comprehensive defect detection system, including both localization and classification of solder joints. For the classification of solder joints, a method integrating active learning and SSL is employed. Initially, an SVM classifier is trained on a small set of labelled data. During the training process, K-means clustering is utilized to analyze the distribution structure of the data, enabling the automatic or manual annotation of a portion of unlabeled samples. This expansion of the labeled dataset contributes to improving the performance of the classifier. Furthermore, (chen2024weakly, ) proposes a weakly SL method leveraging multi-directional integrated circuit markings for surface defect inspection in surface mount technology. This approach employs a combination of weakly SL techniques and integrated circuit recognition to achieve high classification performance with minimal annotation requirements. Furthermore, to enhance defect detection performance, other learning methods are also integrated into SSL approaches. (chu2020neural, ) proposes a semi-supervised framework for anomaly detection and segmentation to identify and segment abnormal regions in images. It employs an autoencoder to generate the reconstruction loss of images and combines it with an RL-based neural batch sampler to provide training batches for the autoencoder, thereby improving the performance of the predictor.
Compared to unsupervised methods, SSL significantly improves model performance when provided with a small amount of labeled defective samples. This approach is beneficial in addressing new types of defects in industrial product manufacturing. The application of SSL is primarily focused on defect classification at the image level, while research on defect localization tasks is still an area that requires further exploration.
@@ 3.4. Reinforcement Learning Algorithms
MDD often faces complex and dynamic environments, where the types and features of defects may exhibit diversity. Reinforcement learning (RL) can learn optimal strategies through interaction with the environment, enhancing the effectiveness of MDD. Therefore, RL, capable of meeting the demands of more complex material defect detection environments and achieving autonomous real-time dynamic learning of material defects, has started to gain favour among researchers.
@@
In (chu2020neural, ), an RL-based neural batch sampler optimizes training batches for an autoencoder, enhancing its ability to distinguish abnormal and normal regions through loss curves. Periodic reinitialization and retraining allow the autoencoder to optimize towards diverse local minima, improving prediction performance. Additionally, (chung2022reinforcement, ) introduces Continuous G-Learning, an RL-based online defect detection method that leverages offline and online prior knowledge to adjust process parameters, ensuring rapid defect mitigation and improved quality assurance in additive manufacturing. The study also proposes an RL-based virtual texture surface defect generation system, where actor-critic networks search for optimal defect rendering parameters, generating realistic images that enhance training data diversity and detection recall rates. Furthermore, (zhang2023duak, ) develops a bidirectional RL-based knowledge graph inference method for steel surface defect detection. This approach employs dual agents to explore paths in opposite directions, leveraging historical experiences to update policies, improving inference accuracy, and supplementing missing knowledge. Lastly, (mohanty2024enhancing, ) integrates a Proximal Policy Optimization (PPO) algorithm within a digital twin model of titanium alloy washers, enabling an RL agent to simulate realistic defects for training a ResNet18-based detector, enhancing its robustness and real-world performance.
@@
RL methods can adapt to continuously changing environments in MDD. This is because they learn through interaction with the environment, enabling them to continuously adjust their strategies during training to accommodate new defect types or environmental changes. They also exhibit high real-time capability, allowing learning and decision-making in real-time data streams, which is crucial for scenarios requiring timely responses in MDD. Additionally, RL models can autonomously learn defect features without the need for extensive manual feature engineering. This enables them to extract valuable information from data and exhibit adaptability to complex defect situations. However, RL typically demands substantial training and computational resources, especially when dealing with complex material defect detection tasks. This may result in long training times and high computational costs. Moreover, RL models usually require a large number of samples for training, particularly in the case of deep RL models. Acquiring large-scale annotated data in the field of MDD can be an expensive and challenging task. Overall, RL holds potential in MDD.
@@ 3.5. Generative Learning Algorithms
To obtain more diverse augmentation data, generative learning (GL) techniques are introduced into the augmentation of defect sample data, aiming to synthetically create realistic defects on normal samples. The work by (valente2020print, ) introduces an end-to-end deep learning framework based on semantic segmentation for mapping printing defects. This framework enables pixel-level detection and localization of defects, such as stripes and colour bands, in printed images. Two types of printing defects and printing scan effects are simulated using image processing and computer graphics techniques, generating synthetic training datasets. This approach avoids the costs and time associated with manually collecting and annotating data. Experimental results demonstrate that the proposed framework can effectively learn features from synthetic data and generalize them to real printing defects. In (bosse2023automated, ), a data-driven feature labeling model is proposed for detecting hidden defects, such as pores in aluminium alloy die-cast plates, in low-cost X-ray projection images. Due to the absence of accurate annotated data for real X-ray images, the authors use numerical simulations of X-ray images to generate synthetic training datasets, taking into account various noise and distortion factors. Furthermore, (farady2023preaugnet, ) introduces an offline PreAugmentation Network (PreAugNet) for generating and filtering augmented samples to improve the training of target networks. The PreAugmentation Network utilizes a CNN to extract features from original samples and updates the generated samples through the decision boundary analysis of an SVM classifier. This ensures that the new samples produced by the PreAugmentation Network are more diverse and representative. Evaluation on three real industrial steel surface defect datasets demonstrates the effectiveness of PreAugmentation Network data augmentation in improving the training of defect classification networks.
@@
In addition to the methods mentioned above for generating defect data, the most common generative techniques are autoencoders and GANs. (yun2020automated, ) proposes a metal surface defect classification algorithm using Deep Convolutional Neural Networks (DCNN) and a Conditional Convolutional VAE (CCVAE) for data augmentation. By learning the distribution of defect data, CCVAE generates diverse defect images, enhancing the DCNN model’s generalization and accuracy. Similarly, GAN-based methods are widely utilized to generate defect images. (pourreza2021g2d, ) introduces G2D (Generate to Detect Anomaly), leveraging GANs to produce anomaly samples during training, enabling anomaly detection in images and videos with only normal samples. (zhang2021defect, ) presents Defect-GAN, which simulates damage and restoration processes to generate realistic defect samples, addressing the scarcity of defect data. Furthermore, (jain2022synthetic, ) employs three GAN architectures (DCGAN, AC-GAN, and InfoGAN) to generate defect images for six types of steel strip defects. Their study explores latent variables in GANs to control synthetic image features like defect severity and lighting, incorporating domain knowledge and improving classification performance. These generative techniques demonstrate significant potential in augmenting datasets, particularly in scenarios with limited real-world defect samples.
@@
In summary, GL methods can be employed in MDD to synthesize defect samples, thereby increasing the diversity of training data. This is particularly beneficial for MDD tasks that lack a large number of annotated defect samples, as it can significantly alleviate the cost of defect annotation, address data imbalance issues, and enhance model robustness. However, the performance of GL methods largely depends on the quality of the generated model, which may be constrained by the limitations of training data and may struggle to fully replicate the real-world distribution of defects. Additionally, synthesized defect samples may not be sufficiently diverse, leading to a decline in model performance in practical applications. Moreover, some GL methods, especially those based on DL, may have high computational resource requirements, especially when training and generating on a large scale. Therefore, the choice of GL methods should be comprehensively considered based on specific task requirements, data characteristics, and computational resources.